OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 20°C and 160°C. The temperature is at its lowest point when t = 0, the temperature is 90 degrees and increasing. What is a sine function that would model this reaction?

OpenStudy (anonymous):

The Answers are: f(t) = 90 sin pi over 4t + 70 f(t) = 70 sin pi over 4t + 90 f(t) = 70 sin 8t + 90 f(t) = 90 sin 8t + 70

OpenStudy (anonymous):

I am thinking its B or C since the Amplitude is 70 and the Vertical Shift is 90.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (aum):

In the second choice, is t in the denominator or the numerator?

OpenStudy (anonymous):

Neither, you are multiplying it to pi/4.

OpenStudy (aum):

\(\Large \sin(\frac{\pi}{4}t)\)?

OpenStudy (anonymous):

\[f(t)=70\sin(\pi/4)t+90\]

OpenStudy (anonymous):

Is the second answer

OpenStudy (aum):

Put t= 0 in each answer choice and evaluate f(t). It should be 90. You can quickly rule out two of the choices.

OpenStudy (anonymous):

???

OpenStudy (aum):

Something doesn't seem right. It says the temperature ranges from 20 to 160. That means the lowest temperature is 20. But then it says: "temperature is at its lowest point when t = 0, the temperature is 90 degrees" implying the lowest temperature is 90 degrees!???

OpenStudy (anonymous):

I don't know, this question is strange.

OpenStudy (aum):

Can you post a screenshot of the question and the answer choices?

OpenStudy (anonymous):

I think its B

OpenStudy (anonymous):

Or maybe C?

OpenStudy (aum):

Both b and c satisfy all the other conditions except the "temperature is at its lowest point when t = 0".

OpenStudy (anonymous):

I will just answer B

OpenStudy (aum):

Try B. If it does not accept that, try C. But I think there may be an error in the question.

OpenStudy (anonymous):

K

OpenStudy (anonymous):

Since your here, can you help with this other question?

OpenStudy (aum):

I can try.

OpenStudy (anonymous):

The graph of the sine curve below is of electromagnetic energy that represents red light:

OpenStudy (anonymous):

What function accurately represents the sine curve for red light?

OpenStudy (aum):

From the graph, the period is 640 nanometers. In f(theta) = Asin(B * theta), the period is 2pi / B Equate 2pi/B to 640 and solve for B.

OpenStudy (anonymous):

The Answers are: A:f(x) = sin pi over 640x B:f(x) = sin 640πx C:f(x) = sin 320πx D:f(x) = sin pi over 320x

OpenStudy (aum):

I can't give answers. Equate 2pi/B to 640 and solve for B. Then put it in f(x) = sin(B * x)

OpenStudy (anonymous):

what was the answer to your original question?

OpenStudy (anonymous):

like was it b or c?

OpenStudy (anonymous):

@Johnny_Boy_14

OpenStudy (anonymous):

???

OpenStudy (anonymous):

It was C

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

okay thank you!